# Proca stars in excited states

### v1

@@ Acknowledgements.
We would like to thank Jose Damian Lopez, Claudio Lazarte, Axel Rangel and Jorge Yahir Mio for many useful discussions and comments. This work was partially supported by CONAHCYT Network Projects No. 376127 and No. 304001, and DGAPA-UNAM project IN100523.
@@ Appendix A Migration data
The last two columns in the Table show the frequency $\omega$ and mass $M$ for those solutions in the ground state toward which the perturbed excited solutions seem to be migrating. It is important to mention the fact that the frequencies of the ground state solutions $\omega$ used for the comparisons are not exactly the same as those of the final stage of the migrating solutions $\omega_{f}$. The reason for this is that we do not have solutions for all values of $\omega$ since we have only sampled the solution space for a finite (though large) number of cases, and the frequency is not our free parameter but rather the eigenvalue found at the end. However, we choose those solutions in the ground state that fall within the uncertainty range of $\omega_{f}$.
With these caveats, once we choose the comparison frequency $\omega$ for a fround state solution the corresponding mass $M$ is fixed, and it is remarkable how close it is to the value $M_{f}$ of the migrating solution in each case. This convinces us that the perturbed excited states are indeed migrating to stable solutions on the ground state.

### v2

@@ Acknowledgements.
We thank Jose Damian Lopez, Claudio Lazarte, Axel Rangel and Jorge Yahir Mio for many useful discussions and comments. This work was partially supported by CONAHCYT Network Projects No. 376127 and No. 304001, and DGAPA-UNAM project IN100523.
@@ Appendix A Migration data
The last two columns in the table show the frequency $\omega$ and mass $M$ for those solutions in the ground state toward which the perturbed excited solutions seem to be migrating. It is important to mention the fact that the frequencies of the ground state solutions $\omega$ used for the comparisons are not exactly the same as those of the final stage of the migrating solutions $\omega_{f}$. The reason for this is that we do not have solutions for all values of $\omega$ since we have only sampled the solution space for a finite (though large) number of cases, and the frequency is not our free parameter, but rather the eigenvalue found at the end. However, we choose those solutions in our family of sampled ground state configurations for which the frequency is closest to $\omega_{f}$.
With these caveats, once we choose the comparison frequency $\omega$ for a ground state solution, the corresponding mass $M$ is fixed, and it is remarkable how close it is to the value $M_{f}$ of the migrating solution in each case. This convinces us that the perturbed excited states are indeed migrating to stable solutions on the ground state.